1. Market demand is given by p = 12 - Q. There are two firms: the incumbent firm (1) and the entrant firm (E). Incumbent moveIf the incumbent can purchase new equipment: 1. he can either produce with the old equipment at marginal cost c =8, or 2 spend K = 5 for new equipment which cust its marginal cost to c = 6.

Questions:

1. If the incumbent had purchased new equipment and anticipates entry, what quantity does he produce?

2. Will he enter the market?

3. What is the incumbent's payoff?

4. Will the incumbent choose to purchase new equipment in the first place?
Answer #1

Market demand, P= 12-0 Case I: When there is only single c=8, firm in comment, do T = PXQ - co = (12-0) Q - 80 ani - 12-20-8=ni= 1291 - 9,² - (491-9; 2) - 891 dą - 12 - 291-(4-291) - 8 = 0 dar 12-291-2 +291-8=0 2 7 Q2 = 4.91 - 4 - 2 = 2 / Pe 12-91-96 he had do he had purchased how equipment, and anticipates enby then (Q = 9 units. 2. Eiherant can proo 3. Perant can producAnswer question